# Zhang, Liu, and Zhao

### v3

@@ 10 Details of Convergence Analysis in Section 4
$\mathcal{E}_{1}^{\nu}$ $=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})-\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
@@
where the last inequality holds by Lemma 10.4, and the Lipschitz continuity of surrogates follows from the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$, compactness of $\Theta$, and finiteness of $S$. Then the second inequality in the proposition can be proved with arguments similar to those above.∎
10.3 Proofs of Theorems 4.4 and 4.5
In this section, we prove the convergence results in Theorems 4.4 and 4.5. As in Section 4.2, we suppose throughout this section that $\varphi$ is B-differentiable, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ satisfies Assumption 1 for each $\xi\in S$.
The convergence analysis requires the following two residual functions. The first residual of a reference point $\theta^{\prime}\in\Theta$ with $\varepsilon\geq 0$ and $\rho>0$ measures the proximal mapping gap with respect to the surrogation function $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime})\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},I)$, which is defined as
$r^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\max_{\theta\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta^{\prime})}\left\lVert\theta-\theta^{\prime}\right\rVert,\text{\ where\ }\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime})+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}.$ (21)
Clearly, $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (9) if and only if $r^{\varepsilon,\rho}(\theta^{d})=0$. The residual $r^{\varepsilon,\rho}$ works as a criterion for the subsequential convergence to composite strong d-stationarity and is also exploited to establish the error-bound analysis. The other useful residual is defined as $r_{d}(\theta)\coloneqq\underset{d\in\mathcal{T}(\theta;\Theta),\left\lVert d\right\rVert\leq 1}{\max}\left\{-F^{\prime}(\theta;d)\right\}$, where $\mathcal{T}(\theta;\Theta)$ is the tangent cone of $\Theta$ at $\theta\in\Theta$. This residual is utilized to derive the global error bound in terms of $r^{\varepsilon,\rho}$ in the second statement of Proposition 10.6. In Figure 10.3, we summarize a roadmap for the convergence analysis of the enhanced SMM algorithm.
A Roadmap for Convergence Analysis of Enhanced SMM.
@@
    Under Assumptions 2 and 1, for $\varepsilon>0$ and $\rho>0$, there exists a constant $\eta_{r}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}r^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$. In particular, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, the error bound holds for any $\varepsilon\geq 0$ and $\rho>0$ without the two assumptions.
Before proving this proposition, we show that Theorem 4.4 could be derived straightforwardly by combining Propositions 4.3 and 10.6, similarly for Theorem 4.5.
Proof 10.7
Proof of Theorem 4.4. According to the definition in (21), for any $\theta\in\mathcal{P}^{\varepsilon,\rho}_{\min}(\theta^{\prime})$, there exists $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ such that $\theta=P^{\rho}_{I}(\theta^{\prime})$. Due to the finite-sum structure of the objective $F$, there exists a uniform constant $C_{\mathcal{I}}>0$ such that the cardinality $\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert\leq C_{\mathcal{I}}$ for any $\theta^{\prime}\in\Theta$, and thus ${r}^{\varepsilon,\rho}(\theta^{\prime})\leq\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert\bar{r}^{\varepsilon,\rho}(\theta^{\prime})\leq C_{\mathcal{I}}\bar{r}^{\varepsilon,\rho}(\theta^{\prime})$, which implies that $\mathbb{E}\,[{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ converges to 0 at least at the same rate as $\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ in terms of $\{N_{\nu}\}_{\nu}$ and $T$. Therefore, the statements of the theorem follow from Propositions 4.3 and 10.6.∎
We next present the proof of Proposition 10.6. The overall scheme of the error-bound analysis follows the error-bound theory developed in Liu et al. (2022) for the post-convergence analysis of the SMM algorithm. To justify the global error bound in the second statement of Proposition 10.6, we first give the error bounds of distance to $\mathcal{S}^{d}$ in terms of the residual $r_{d}$ in Proposition 10.8 and then relate $r_{d}$ with $r^{\varepsilon,\rho}$ in Proposition 10.10.
@@
For problem (9), the following statements hold:
  1. (a)
    If $\varphi$ is piecewise affine and $\Theta$ is a polyhedral set, then there exists $\eta_{d}>0$ such that
@@
  2. (b)

### v4

@@ 7 Surrogation Construction of $F(\cdot,\xi)$ for Case 1 to Case 3
$\mathcal{E}_{1}^{\nu}$ $=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})\,\middle|\,\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})-\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
@@
where the last inequality holds by Lemma 8.4, and the Lipschitz continuity of surrogates follows from the bi-variate continuity of $\widehat{F}(\cdot,\xi;\cdot,I)$, compactness of $\Theta$, and finiteness of $\Xi$. Then the second inequality in the proposition can be proved with arguments similar to those above.∎
8.3 Proofs of Theorems 3.6 and 3.7
In this section, we prove the convergence results in Theorems 3.6 and 3.7. As in Section 3.2.3, we suppose throughout this section that $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}$ of $F(\cdot,\xi)$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$.
The convergence analysis requires the following two residual functions. The first residual of a reference point $\bar{\theta}\in\Theta$ with $\varepsilon\geq 0$ and $\rho>0$ measures the proximal mapping gap with respect to the surrogate function $\widehat{F}^{\varepsilon}_{\min}(\theta;\bar{\theta})\coloneqq\min_{\mathbf{I}^{\varepsilon}\in\mathcal{I}^{\varepsilon}(\bar{\theta})}\widehat{F}(\theta;\bar{\theta},\mathbf{I}^{\varepsilon})$, which is defined as
$r^{\varepsilon,\rho}(\bar{\theta})\coloneqq\max_{\theta\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\bar{\theta})}\left\lVert\theta-\bar{\theta}\right\rVert,\text{\ where\ }\mathcal{P}_{\min}^{\varepsilon,\rho}(\bar{\theta})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta;\bar{\theta})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}.$ (29)
Clearly, $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (10) if and only if $r^{\varepsilon,\rho}(\theta^{d})=0$. The residual $r^{\varepsilon,\rho}$ works as a criterion for the subsequential convergence to composite strong d-stationarity and is also exploited to establish the error-bound analysis. The other useful residual is defined as $r_{d}(\theta)\coloneqq\underset{d\in\mathcal{T}(\theta;\Theta),\left\lVert d\right\rVert\leq 1}{\max}\left\{-F^{\prime}(\theta;d)\right\}$, where $\mathcal{T}(\theta;\Theta)$ is the tangent cone of $\Theta$ at $\theta\in\Theta$. This residual is utilized to derive the global error bound in terms of $r^{\varepsilon,\rho}$ in the second statement of Proposition 8.6. In Figure 8.3, we summarize a roadmap for the convergence analysis of the ESMM algorithm.
[…]
A roadmap for convergence analysis of ESMM
@@
    For $\varepsilon\geq 0$ and $\rho>0$, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, there exists a constant $\eta_{r}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}r^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$.
Before proving this proposition, we show that Theorem 3.6 could be derived straightforwardly by combining Propositions 3.5 and 8.6, similarly for Theorem 3.7.
Proof 8.7
Proof of Theorem 3.6. According to the definition in (29), for any $\theta\in\mathcal{P}^{\varepsilon,\rho}_{\min}(\bar{\theta})$, there exists $\mathbf{I}^{\varepsilon}\in\mathcal{I}^{\varepsilon}(\bar{\theta})$ such that $\theta=P^{\rho}_{\mathbf{I}^{\varepsilon}}(\bar{\theta})$. Due to the finite-sum structure of the objective $F$, there exists a uniform constant $C_{\mathcal{I}}>0$ such that the cardinality $\left\lvert\mathcal{I}^{\varepsilon}(\bar{\theta})\right\rvert\leq C_{\mathcal{I}}$ for any $\bar{\theta}\in\Theta$, and thus ${r}^{\varepsilon,\rho}(\bar{\theta})\leq\left\lvert\mathcal{I}^{\varepsilon}(\bar{\theta})\right\rvert\bar{r}^{\varepsilon,\rho}(\bar{\theta})\leq C_{\mathcal{I}}\bar{r}^{\varepsilon,\rho}(\bar{\theta})$, which implies that $\mathbb{E}\,[{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ converges to 0 at least at the same rate as $\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ in terms of $\{N_{\nu}\}_{\nu}$ and $T$. Therefore, the statements of the theorem follow from Propositions 3.5 and 8.6.∎
[…]
@@
For problem (10), if $\varphi$ is piecewise affine and $\Theta$ is a polyhedral set, then there exists $\eta_{d}>0$ such that
@@
Proof 8.9
Proof of Proposition 8.8. Since the objective $F$ is a piecewise affine function on a polyhedral set $\Theta$, $\Theta$ can be decomposed into finitely many polyhedral sets such that $F$ is affine on each of them, showing that $r_{d}(\cdot)$ takes finitely many values on $\Theta$. Notice that $\theta\in\mathcal{S}^{d}$ if and only if $r_{d}(\theta)=\max_{d\in\mathcal{T}(\theta;\Theta),\left\lVert d\right\rVert\leq 1}\{-F^{\prime}(\theta;d)\}=0$. Hence, there exists $l_{0}>0$ such that $r_{d}(\theta)\geq l_{0}$ for any $\theta\in\Theta\backslash\mathcal{S}^{d}$. By setting $\eta_{d}\coloneqq\frac{2}{l_{0}}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, we deduce $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq 2\max_{\theta\in\Theta}\left\lVert\theta\right\rVert\leq\eta_{d}r_{d}(\theta)$. ∎
Proposition 8.10
Suppose that the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies the quadratic surrogation gap condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$. For problem (10) with $\varepsilon\geq 0$ and $\rho>0$, it holds that for any $\theta\in\Theta$, there exists $\theta^{\prime}\in\Theta$ such that $\left\lVert\theta^{\prime}-\theta\right\rVert\leq 2\,r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\theta^{\prime})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}(\theta)$.
[…]
Proof 8.11
Proof of Proposition 8.6. For the first statement, let $\{\theta_{t}\}_{t\in\mathcal{K}}$ be a convergent subsequence with the limit point $\theta_{\infty}$. Given $\varepsilon^{\prime}\in[0,\varepsilon)$, it follows from the Lipschitz continuity of $f$ that $\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\subseteq\mathcal{I}^{\varepsilon}(\theta_{t})$ for sufficiently large $t\in\mathcal{K}$. With $\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})\coloneqq\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta_{t},\mathbf{I})+\frac{\rho}{2}\left\lVert\theta-\theta_{t}\right\rVert^{2}\right\}$, by definition of the residual $r^{\varepsilon,\rho}$ there exists $\mathbf{I}^{*}_{t}\in\underset{\mathbf{I}\in\mathcal{I}^{\varepsilon}(\theta_{t})}{\arg\min}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t});\theta_{t},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})-\theta_{t}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon,\rho}(\theta_{t})=\,\big{\lVert}\,{\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})-\theta_{t}}\,\big{\rVert}$. Then for sufficiently large $t\in\mathcal{K}$, we have that
$\widehat{F}^{*}(\theta_{t})$ $\coloneqq\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t});\theta_{t},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
$\geq\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon}(\theta_{t})}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t});\theta_{t},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
$=\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t});\theta_{t},\mathbf{I}^{*}_{t})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}$
$\geq F(\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t}))+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}.$ (30)
[…]
$F(\theta_{\infty})\leq\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{F}^{*}(\theta_{t})=\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{\infty});\theta_{\infty},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\},$
which leads to $F(\theta_{\infty})<\displaystyle{\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\backslash\mathcal{I}^{0}(\theta_{\infty})}}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\theta_{\infty});\theta_{\infty},\mathbf{I})+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\}$ for $\rho^{\prime}>\rho$, yielding $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$.
We next prove the second statement. If $\varphi$ is piecewise affine, by Proposition 3.2 in Appendix 7, the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies Condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$. Given $\varepsilon\geq 0$ and $\rho>0$, by Proposition 8.8 there exists $\eta_{d}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{d}r_{d}(\theta)$ for any $\theta\in\Theta$; by Proposition 8.10, for any $\theta\in\Theta$, there exists $\theta^{\prime}\in\Theta$ such that $\left\lVert\theta^{\prime}-\theta\right\rVert\leq 2r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\theta^{\prime})\leq(2\rho+\frac{5\kappa}{2})r^{\varepsilon,\rho}(\theta)$. Setting $\eta_{r}=2+\eta_{d}\left(2\rho+\frac{5\kappa}{2}\right)$, we have that
$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})+\left\lVert\theta-\theta^{\prime}\right\rVert\leq\eta_{d}r_{d}(\theta^{\prime})+2r^{\varepsilon,\rho}(\theta)\leq\eta_{r}r^{\varepsilon,\rho}(\theta),$
which completes the proof.∎
8.4 Nonasymptotic convergence result for general cases
In this section, we extend the nonasymptotic convergence result for the general case where the outer function $\varphi$ is directionally differentiable and locally Lipschitz continuous, and the surrogate functions of $F$ satisfy Assumptions (B1)-(B4). The error-bound analysis in this case requires the following two assumptions. We denote the closure operator of a set by $\mathrm{cl}(\cdot)$ and $\mathcal{P}_{\min}^{\varepsilon,\rho}(\bar{\theta})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta;\bar{\theta})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}$.
  1. (B5)
    The minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies the quadratic surrogation gap condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$;
  2. (B6)
Assumption (B5) has appeared in Proposition 8.10, and we state it here officially. Condition (P4) originating from Liu et al. (2022) is a technical condition for the nonasymptotic convergence analysis of the ESMM algorithm. Note that Condition (P4) is verified for Case 1-Case 3 in Online Appendix 7. Assumption (B6) is generalized from Drusvyatskiy and Lewis (2018). The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Online Appendix 8.5. It can be expected to hold in more general cases since it is a local property only requiring the convexity and quadratic growth of $F$ near $\mathcal{S}^{d}$.
With those assumptions, the nonasymptotic convergence result of ESMM for general cases is stated as follows.
Theorem 8.12 (Nonasymptotic convergence of ESMM for general cases)
Under Assumptions (B1)-(B6), for problem (10) under the same settings of Algorithm 1 specified in Theorem 3.6, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$.
Following the error-bound analysis developed in Section 8.3, the convergence result can be established with the following proposition.
Proposition 8.13
Under Assumptions (B1)-(B6), for $\varepsilon>0$ and $\rho>0$, there exists a constant $\eta_{r}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}r^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$.
Similar to the proof of the second statement of Proposition 8.6, to prove the above error bound, we require the following proposition, which is a generalization of Proposition 8.8.
Proposition 8.14